# Curbing PBHs with PTAs

### v1

@@ III PBH formation
$\frac{{\rm d}N_{k}}{{\rm d}\nu}=\frac{4\pi}{3}r_{m}^{3}\mathcal{N}$ (17)
and, since $N_{k}\ll 1$, the probability of finding at least one peak within a range $(\nu,\nu+{\rm d}\nu)$ within a Hubble volume is $P_{k}\approx{\rm d}N_{k}$. The mass fraction of PBHs is then Yoo et al. (2018, 2019); Gow et al. (2021); Franciolini et al. (2022)
$P_{k}(\mathcal{C})=\frac{1}{3\pi}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)$ (18)
and the domain of integration agrees with Eq. 8. Thus, peaks theory predicts an additional $\nu^{3}$ factor to the collapse probability in the Gaussian case.
Recasting the quantities in terms of $M_{k}$, as in (III), we can express the PBH abundance as
@@
×$\times\frac{{\cal K}}{\gamma}\frac{\left(2\Phi\left(1-\sqrt{\Lambda}\right)\right)^{3}}{3\pi\sigma_{c}^{4}\Lambda^{1/2}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}$
×$\times\exp\left[{\frac{-2\Phi^{2}}{\sigma_{c}^{2}}\left(1-\sqrt{\Lambda}\right)^{2}}\right]\,.$ (19)
IV Scalar-induced gravitational waves
Primordial curvature fluctuations large enough to generate PBHs will also induce a non-negligible SIGW background. At the second order in perturbation theory, scalar modes can source tensor modes (for a review, see Ref. Domènech (2021)). Starting from Einstein’s equations and omitting anisotropic stress, the equations of motion for GWs in Fourier space are
$h^{\prime\prime}_{\lambda}(\textbf{k},\eta)+2\mathcal{H}h^{\prime}_{\lambda}(\textbf{k},\eta)+k^{2}h_{\lambda}(\textbf{k},\eta)=4\mathcal{S}_{\lambda}(\textbf{k},\eta)\,,$ (20)
with $\lambda$ indicating the polarization and the being the derivative with respect to the conformal time $\eta$. The source term on the right-hand side reads Baumann et al. (2007)
$\mathcal{S}_{\lambda}(\textbf{k},\eta)\!=\!\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})f(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (21)
where $Q$ is a projection factor and $f$ is a function that contains the linear evolution of the Newtonian potential after horizon re-entry and the transfer function mapping them to the curvature perturbation $\zeta$. Eq. (20) can be solved using the Green’s functions approach
$h_{\lambda}(\textbf{k})\!=\!4\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})I(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (22)
where $I$ is the transfer function. To speed up our analysis, we assume perfect radiation domination and do not account for the variation of sound speed during the QCD era (see, for example, Hajkarim and Schaffner-Bielich (2020); Abe et al. (2021)) which also leads to specific imprints in the low-frequency tail of any cosmological SGWB Franciolini et al. (2024). On top of that, cosmic expansion may additionally be affected by unknown physics in the dark sector, which can, e.g., lead to a brief period of matter domination of kination Ferreira and Joyce (1998); Pallis (2006); Redmond et al. (2018); Co et al. (2022); Gouttenoire et al. (2021); Chang and Cui (2022). Explicit analytical expressions for the kernel in different epochs can be found in Domènech (2020), after performing the oscillation average well inside the horizon.
@@ IV Scalar-induced gravitational waves
$\langle h_{\lambda_{1}}(\mathbf{k}_{1})h_{\lambda_{2}}(\mathbf{k}_{2})\rangle=\delta^{3}(\mathbf{k}_{1}+\mathbf{k}_{2})\delta_{\lambda_{1}\lambda_{2}}\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{h,\lambda_{1}}(k_{1})$ (23)
@@
$\Omega_{\rm GW}(k,\eta)=\frac{1}{48}\left(\frac{k}{a(\eta)H(\eta)}\right)\sum_{\lambda=+,\times}\overline{\mathcal{P}_{h,\lambda}(k)}$ (24)
@@
$\Omega_{\rm GW}(k,\eta)|_{\rm G}=$ $\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}\int_{0}^{\infty}{\rm{d}}t\int_{-1}^{1}{\rm{d}}s\hskip 2.84544pt\frac{1}{u^{2}v^{2}}$
×$\times\overline{\tilde{J}^{2}(u,v,x)}\hskip 2.84544pt\mathcal{P}_{\zeta_{g}}\mathcal{P}_{\zeta_{g}}(vk)\mathcal{P}_{\zeta_{g}}(uk)\,,$ (25)
@@
$h^{2}\Omega_{\rm GW}|_{k,0}=h^{2}\Omega_{\rm rad,0}\left(\frac{g_{s,\eta_{0}}}{g_{s,\eta_{k}}}\right)^{\frac{4}{3}}\frac{g_{\rho,\eta_{k}}}{g_{\rho,\eta_{0}}}\Omega_{\rm GW}(k,\eta)\,,$ (26)
@@
where we used the energy ($g_{\rho}$) and entropy density ($g_{s}$) degrees of freedom of Ref. Borsanyi et al. (2016). According to Ref. Aghanim et al. (2020), we consider $h^{2}\Omega_{\rm rad,0}=4.2\cdot 10^{-5}$.
IV.1 Effects of primordial NG on SIGW
@@
As shown in Kohri and Terada (2018); Cai et al. (2019a, b); Hajkarim and Schaffner-Bielich (2020); Atal and Domènech (2021); Yuan and Huang (2021); Domènech et al. (2022); Garcia-Saenz et al. (2023); Liu et al. (2024); Yuan et al. (2023); Li et al. (2023); Unal (2019); Adshead et al. (2021); Abe et al. (2023); Perna et al. (2024) primordial NGs can leave a non-negligible imprint on the SIGW power spectrum. Hence, considering only the quadratic contribution $f_{\rm NL}$ in the local expansion in Eq. (5) [^3] We specify that, in principle, all the terms in the power series expansion in Eq. (5) should be included. Being interested mainly in the qualitative effect of NGs, we truncated the power series at the quadratic contribution, $f_{\rm NL}$, setting all other terms to 0. These additional NG corrections would further shift the results in Fig. 3., at the next-to-leading order one obtains (see, e.g., Unal (2019); Adshead et al. (2021))
$\Omega_{\rm GW}(k,\eta)|_{\rm t}$ $=\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$
$\times$ $\cos 2\varphi_{12}\frac{u_{1}v_{1}}{(u_{2}v_{2})^{2}}\frac{1}{w_{a,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(u_{2}k)\mathcal{P}_{\zeta_{g}}(w_{a,12}k)\,,$ (27)
$\Omega_{\rm GW}(k,\eta)|_{\rm u}$ $=\hskip 2.84544pt\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$
$\times$ $\cos 2\varphi_{12}\frac{u_{1}u_{2}}{(v_{1}v_{2})^{2}}\frac{1}{w_{b,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(v_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(w_{b,12}k)$ (28)
$\Omega_{\rm GW}(k,\eta)|_{\rm hyb}$ $=\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}$
$\times$ $\frac{1}{(u_{1}v_{1}u_{2}v_{2})^{2}}\hskip 2.84544pt\overline{\tilde{J}^{2}(u_{1},v_{1},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(u_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}v_{1}k)\mathcal{P}_{\zeta_{g}}(u_{2}v_{1}k)\,.$ (29)
@@
As in the Gaussian case, the $u_{i}$ and $v_{i}$ are related similarly to the integration variables $t_{i}$ and $s_{i}$.
The GW spectrum can then be expressed as
$\Omega_{{\rm GW}}|_{k,0}$ $=A^{2}\Omega_{\rm GW}|_{{\rm G}}(k/k_{*},0)$ (30)
$+A^{3}f_{\rm NL}^{2}\Omega_{\rm GW}|_{{\rm NG}}(k/k_{*},0)\,,$
@@
Figure 1: Left panel: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum (1) with $\Delta=0.5$ assuming Gaussian fluctuations and changing the window function. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, assuming NGs arising solely from the non-linear corrections. Right panel: Inferred constraints on PBH abundance.
V Constraints on PBHs
V.1 PTA observations
Consider the constraints on the SIGW signal from the current PTA observations. Using from the NANOGrav 15 year dataset Agazie et al. (2023a); Collaboration (2023), we compute the observational upper bound at 95% confidence level on $h^{2}\Omega^{\rm bound}_{\rm GW}$ in each of the first 14 frequency bins. The frequency of bin ”$i$” is given by $f_{i}=i/T$, where $T=16.03\,{\rm yr}=(1.98\,{\rm nHz})^{-1}$ is the time of observation. We then compare the observation with theoretical prediction $h^{2}\Omega_{\rm GW}$ in Eq. (30) in each frequency bin. The constraint on $h^{2}\Omega_{\rm GW}$ in any given bin implies an upper bound on $A$ for a given model of NGs and a fixed shape of the curvature power spectrum. To combine the constraints of individual bins, we consider the strongest constraint on $A$. The upper bound on $A$ is then mapped to an upper bound on $f_{\rm PBH}$. More specifically, for every fixed benchmark model in Table 1, we can map the pair $(A,k_{*})$ uniquely into the pair $(f_{\rm PBH},\langle M_{\rm PBH}\rangle)$.
Due to the finite observational time, the spectral resolution of the detector is limited and can be estimated to be approximately $\Delta f=1/T=1.98\,{\rm nHz}$. This will limit the PTA sensitivity to sharp features in the SIGW spectrum. To account for this, the contribution to $i$-th bin can be estimated as
$\Omega_{{\rm GW},i}=\int\frac{{\rm d}f}{f}\Omega_{\rm GW}^{\rm th}(f)W(f-f_{i})\,$ (31)

### v2

@@ 3 PBH formation
$\frac{{\rm d}N_{k}}{{\rm d}\nu}=\frac{4\pi}{3}r_{m}^{3}\mathcal{N}$ (3.12)
and, since $N_{k}\ll 1$, the probability of finding at least one peak within a range $(\nu,\nu+{\rm d}\nu)$ within a Hubble volume is $P_{k}\approx{\rm d}N_{k}$. The mass fraction of PBHs is then [134, 135, 136, 137]
$P_{k}(\mathcal{C})=\frac{1}{3\pi}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)$ (3.13)
and the domain of integration agrees with Eq. 3.3. Thus, peaks theory predicts an additional $\nu^{3}$ factor to the collapse probability in the Gaussian case.
Recasting the quantities in terms of $M_{k}$, as in (3), we can express the PBH abundance as
@@
×$\times\frac{{\cal K}}{\gamma}\frac{\left(2\Phi\left(1-\sqrt{\Lambda}\right)\right)^{3}}{3\pi\sigma_{c}^{4}\Lambda^{1/2}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\exp\left[{\frac{-2\Phi^{2}}{\sigma_{c}^{2}}\left(1-\sqrt{\Lambda}\right)^{2}}\right]\,.$ (3.14)
4 Scalar-induced gravitational waves
Primordial curvature fluctuations large enough to generate PBHs will also induce a non-negligible SIGW background. At the second order in perturbation theory, scalar modes can source tensor modes (for a review, see Ref. [15]). Starting from Einstein’s equations and omitting anisotropic stress, the equations of motion for GWs in Fourier space are
$h^{\prime\prime}_{\lambda}(\textbf{k},\eta)+2\mathcal{H}h^{\prime}_{\lambda}(\textbf{k},\eta)+k^{2}h_{\lambda}(\textbf{k},\eta)=4\mathcal{S}_{\lambda}(\textbf{k},\eta)\,,$ (4.1)
with $\lambda$ indicating the polarization and the being the derivative with respect to the conformal time $\eta$. The source term on the right-hand side reads [13]
$\mathcal{S}_{\lambda}(\textbf{k},\eta)\!=\!\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})f(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (4.2)
where $Q$ is a projection factor and $f$ is a function that contains the linear evolution of the Newtonian potential after horizon re-entry and the transfer function mapping them to the curvature perturbation $\zeta$. Eq. (4.1) can be solved using the Green’s functions approach
$h_{\lambda}(\textbf{k})\!=\!4\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})I(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (4.3)
where $I$ is the transfer function. To speed up our analysis, we assume perfect radiation domination and do not account for the variation of sound speed during the QCD era (see, for example, [138, 139]) which also leads to specific imprints in the low-frequency tail of any cosmological SGWB [140]. On top of that, cosmic expansion may additionally be affected by unknown physics in the dark sector, which can, e.g., lead to a brief period of matter domination of kination [141, 142, 143, 144, 145, 146, 147, 148, 149]. Explicit analytical expressions for the kernel in different epochs can be found in [150], after performing the oscillation average well inside the horizon.
@@ 4 Scalar-induced gravitational waves
$\langle h_{\lambda_{1}}(\mathbf{k}_{1})h_{\lambda_{2}}(\mathbf{k}_{2})\rangle=\delta^{3}(\mathbf{k}_{1}+\mathbf{k}_{2})\delta_{\lambda_{1}\lambda_{2}}\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{h,\lambda_{1}}(k_{1})$ (4.4)
@@
$\Omega_{\rm GW}(k,\eta)=\frac{1}{48}\left(\frac{k}{a(\eta)H(\eta)}\right)\sum_{\lambda=+,\times}\overline{\mathcal{P}_{h,\lambda}(k)}$ (4.5)
@@
$\Omega_{\rm GW}(k,\eta)|_{\rm G}=$ $\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}\int_{0}^{\infty}{\rm{d}}t\int_{-1}^{1}{\rm{d}}s\hskip 2.84544pt\frac{1}{u^{2}v^{2}}\overline{\tilde{J}^{2}(u,v,x)}\hskip 2.84544pt\mathcal{P}_{\zeta_{g}}(vk)\mathcal{P}_{\zeta_{g}}(uk)\,,$ (4.6)
@@
$h^{2}\Omega_{\rm GW}|_{k,0}=h^{2}\Omega_{\rm rad,0}\left(\frac{g_{s,\eta_{0}}}{g_{s,\eta_{k}}}\right)^{\frac{4}{3}}\frac{g_{\rho,\eta_{k}}}{g_{\rho,\eta_{0}}}\Omega_{\rm GW}(k,\eta)\,,$ (4.7)
@@
where we used the energy ($g_{\rho}$) and entropy density ($g_{s}$) degrees of freedom of Ref. [151]. According to Ref. [152], we consider $h^{2}\Omega_{\rm rad,0}=4.2\cdot 10^{-5}$.
4.1 Effects of primordial NG on SIGW
@@
As shown in [153, 27, 28, 138, 154, 155, 156, 22, 157, 158, 159, 23, 24, 25, 26] primordial NGs can leave a non-negligible imprint on the SIGW power spectrum. Hence, considering only the quadratic contribution $f_{\rm NL}$ in the local expansion in Eq. (2.5) [^4] We specify that, in principle, all the terms in the power series expansion in Eq. (2.5) should be included. Being interested mainly in the qualitative effect of NGs, we truncated the power series at the quadratic contribution, $f_{\rm NL}$, setting all other terms to 0. These additional NG corrections would further shift the results in Fig. 4., at the next-to-leading order one obtains (see, e.g., [23, 24])
$\Omega_{\rm GW}(k,\eta)|_{\rm t}$ $=\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$ (4.8)
$\times$ $\cos 2\varphi_{12}\frac{u_{1}v_{1}}{(u_{2}v_{2})^{2}}\frac{1}{w_{a,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(u_{2}k)\mathcal{P}_{\zeta_{g}}(w_{a,12}k)\,,$
$\Omega_{\rm GW}(k,\eta)|_{\rm u}$ $=\hskip 2.84544pt\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$ (4.9)
$\times$ $\cos 2\varphi_{12}\frac{u_{1}u_{2}}{(v_{1}v_{2})^{2}}\frac{1}{w_{b,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}\mathcal{P}_{\zeta_{g}}(v_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(w_{b,12}k)$
$\Omega_{\rm GW}(k,\eta)|_{\rm hyb}$ $=\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}$ (4.10)
$\times$ $\frac{1}{(u_{1}v_{1}u_{2}v_{2})^{2}}\hskip 2.84544pt\overline{\tilde{J}^{2}(u_{1},v_{1},x)}\mathcal{P}_{\zeta_{g}}(u_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}v_{1}k)\mathcal{P}_{\zeta_{g}}(u_{2}v_{1}k)\,.$
@@
As in the Gaussian case, the $u_{i}$ and $v_{i}$ are related similarly to the integration variables $t_{i}$ and $s_{i}$. The GW spectrum can then be expressed as
$\Omega_{{\rm GW}}|_{k,0}$ $=A^{2}\Omega_{\rm GW}|_{{\rm G}}(k/k_{*},0)+A^{3}f_{\rm NL}^{2}\Omega_{\rm GW}|_{{\rm NG}}(k/k_{*},0)\,,$ (4.11)
@@
Figure 2: Left panel: Constraints from NANOGrav15 [29] on the amplitude of a log-normal curvature power spectrum (2.1) with $\Delta=0.5$ assuming Gaussian fluctuations and changing the window function. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, assuming NGs arising solely from the non-linear corrections. Right panel: Inferred constraints on PBH abundance.
5 Constraints on PBHs
5.1 PTA observations
Consider the constraints on the SIGW signal from the current PTA observations. Using from the NANOGrav 15 year dataset [29, 160], we compute the observational upper bound at 95% confidence level on $h^{2}\Omega^{\rm bound}_{\rm GW}$ in each of the first 14 frequency bins. The frequency of bin ”$i$” is given by $f_{i}=i/T$, where $T=16.03\,{\rm yr}=(1.98\,{\rm nHz})^{-1}$ is the time of observation. We then compare the observation with theoretical prediction $h^{2}\Omega_{\rm GW}$ in Eq. (4.11) in each frequency bin. The constraint on $h^{2}\Omega_{\rm GW}$ in any given bin implies an upper bound on $A$ for a given model of NGs and a fixed shape of the curvature power spectrum. To combine the constraints of individual bins, we consider the strongest constraint on $A$. The upper bound on $A$ is then mapped to an upper bound on $f_{\rm PBH}$. More specifically, for every fixed benchmark model in Table 1, we can map the pair $(A,k_{*})$ uniquely into the pair $(f_{\rm PBH},\langle M_{\rm PBH}\rangle)$.
Due to the finite observational time, the spectral resolution of the detector is limited and can be estimated to be approximately $\Delta f=1/T=1.98\,{\rm nHz}$. This will limit the PTA sensitivity to sharp features in the SIGW spectrum. To account for this, the contribution to the $i$-th bin can be estimated as
$\Omega_{{\rm GW},i}=\int\frac{{\rm d}f}{f}\Omega_{\rm GW}^{\rm th}(f)W(f-f_{i})\,$ (5.1)